OpenStudy (lgbasallote):

create an equation using the most used numbers in mathematics, the most used mathematical symbols, the most used mathematical constants, and the most used greek symbol in mathematics.

OpenStudy (lgbasallote):

Hint: this is a real equation

OpenStudy (anonymous):

most used.. u have to know which is used most.. i assume

OpenStudy (lgbasallote):

hmmm..i guess so :DD

OpenStudy (anonymous):

well, then google it: most used number, most used greek symbol.. blah blah blah

OpenStudy (lgbasallote):

since it's my raison d'etre i'll tag fool @FoolForMath hahaha =)) it's a riddle arnab ;D

OpenStudy (anonymous):

nice question @Igbassalote

OpenStudy (kropot72):

It doesn't meet the spec but this is worth pondering on: \[e ^{i pi}=-1\] Note: The pi symbol is not available right now.

OpenStudy (agreene):

@kropot72 \[\large e^{i\pi}=-1\]

OpenStudy (anonymous):

\[\left\{\lim_{z\to\infty}\frac{\pi\epsilon e!\phi^{-1}\int\tau\sum\delta^{\zeta(\alpha)}\,d\theta\psi_\chi\arg\left(\mod\Omega/\{\omicron\}\right)\gamma K}{\zeta(\alpha)\frac{\partial}{\partial\kappa}\Gamma\left(\int_K^\sigma e^{-t^2}\right)}=0\mid\forall\left(z=\zeta(\alpha),z\in\mathbb Z\right)\right\}\] There! An equation that doesn't mean anything, but has everything I can think of and is technically correct.

OpenStudy (anonymous):

And by "everything I can think of", I mean your question is vague enough to allow me to post something overdone and dumb like this, and have it be technically correct.

OpenStudy (anonymous):

And I fulfill the requirement of having "created" the equation. :|

OpenStudy (karatechopper):

yea but u forgot the fundamentals!! the +, -, /, *.....

OpenStudy (anonymous):

Pfft, fundamentals are for dorks. REAL mathematicians write large, complex equations that mean absolutely nothing. Just kidding. But, no really, the question is a bit vague.

OpenStudy (karatechopper):

loool

OpenStudy (apoorvk):

@BADREFERENCES HAIL HO!!!

OpenStudy (lgbasallote):

@kropot72 got it...i dunno what @badreferences equations meant :p